OpenStudy (theloshua):

|x - 2| = - 1/2 solve

OpenStudy (texaschic101):

| x - 2 | = -1/2 you will have to solve 2 equations for this... x - 2 = -1/2 and -(x - 2) = -1/2 now....can you solve both of those equations for x ?

OpenStudy (theloshua):

x = 1.5 and x = 2.5, right?

OpenStudy (theloshua):

help???

OpenStudy (anonymous):

wouldn't there be no solutions. (Absolute value cannot be less than 0.)

OpenStudy (theloshua):

no, there would be two solutions. thats how you solve absolute value equations. I know that much

OpenStudy (anonymous):

Ok well the you are right on both of the problems

OpenStudy (theloshua):

you sure? :/

OpenStudy (anonymous):

3/2 and 5/2

OpenStudy (anonymous):

Yes i am positive look

OpenStudy (theloshua):

okay :)

OpenStudy (playgames1122):

PEPE THE FROG :D

OpenStudy (anonymous):

Well first of all 3/2 as a decimal is 1.5 and 5/2 is 2.5